The area of the triangle is 36 square inches. The base is half the height. Find the height and base of the triangle.
Found 2 solutions by ewatrrr, stanbon:
Answer by ewatrrr(24785)

 
Hi,
The base is half the height
Let x and 2x represent the base and the height
Question states***
A = 36in^2 = (1/2)x(2x)
36in^2 = x^2
6in = x, the base. (tossing out negative solution for # of inches)
Length is 12in. (2*6in)
CHECKING our Answer***
(1/2)6in*12in = 36in^2

Answer by stanbon(75887)

The area of the triangle is 36 square inches. The base is half the height. Find the height and base of the triangle.
---
Let height = 2x inches
Then base = x inches
---
Area = (1/2)(base)(height)
36 = (1/2)x(2x)
36 = x^2
---
x = 6 inches (base)
2x = 12 inches (height)
===========================
Cheers,
Stan H.
===========================